# Can Large Language Models Understand Symbolic Graphics Programs?

### v1

@@ Abstract
Assessing the capabilities of large language models (LLMs) is important to be able to characterize what they know (and don’t) and how they can be appropriately used. Yet, capability assessments are often challenging, in part, because it is hard to find tasks to which they have not been exposed during training. We take one step to address this challenge by turning to a new task: focusing in on symbolic graphics programs, which are a popular representation for graphics content that procedurally generates visual data. Large language models (LLMs) have shown exciting promise towards program synthesis, but do they “understand” symbolic graphics programs? Unlike conventional programs, symbolic graphics programs can be translated to graphics content (e.g., 2D images, 3D geometry). Here, we characterize an LLM’s “understanding” of symbolic programs in terms of their ability to answer questions related to the graphics (spatial) content. This task is challenging as the questions are difficult to answer from the symbolic programs alone – yet, they would be easy to answer from the corresponding graphics content as we verify through a human experiment. To understand symbolic programs, LLMs may need to possess the ability to “imagine” how the corresponding graphics content would look without directly accessing the rendered visual content. We use this task to evaluate LLMs by creating a large benchmark for the semantic understanding of symbolic graphics programs. This benchmark is built via a novel usage of program-graphics correspondence, hence requiring minimal human efforts. We evaluate both commercial and open-source LLMs on our benchmark to elucidate a preliminary assessment of their ability to reason about visual scenes from programs. We find that this task well distinguishes existing LLMs and models that are considered good at reasoning perform better. Lastly, we introduce a way to improve this ability – Symbolic Instruction Tuning (SIT). Specifically, we query powerful vision-language models (e.g., GPT-4o) with questions and images generated by symbolic programs. These program-question pairs are collected as our instruction dataset which is then used to finetune an LLM. With a small amount of data, we find that SIT can improve the understanding of LLMs regarding symbolic graphics programs. Assessing how well models “understand” symbolic graphics programs offers new possibilities for LLMs to perform visual reasoning. Finally, we showcase such possibilities in generic instruction tuning.
@@ 1 Introduction
What are large language models (LLMs) capable of? Recent studies [4, 56] have shown that LLMs are able to generate generic computer programs, indicating a certain level of understanding about the symbolic structure of programs. Motivated by such progress, we focus on another important family of computer programs, called symbolic graphics programs, where a graphics content (e.g., image, 3D asset) can be generated by running a program. We are interested in the following question: Can large language models “understand” symbolic graphics programs?
Before trying to answer this question, we start by defining what we consider “understanding” of symbolic graphics programs, in the context of this work. Because a (deterministic) graphics program can be uniquely rendered to an image (the graphics programs we consider here), we characterize LLMs’ understanding of the graphics program as the semantic understanding of the corresponding rendered image. More specifically, we approximate such a semantic visual understanding by the ability to correctly answer semantic questions only based on the raw program input. These semantic questions are generated based on the rendered image, such that they are easy to answer given the image and yet challenging given only the program as text prompt. Guided by this insight, we propose a generic pipeline for creating benchmarks that can evaluate this particular ability of LLMs to understand symbolic graphics programs, while requiring minimal human efforts. While we of course recognize that there are other elements of visual reasoning that characterize understanding in humans, and that we want to evaluate in machine intelligences, we believe that our benchmark provides insight into one element of “understanding” over symbolic graphics programs that enables deeper understanding of what current LLMs are (and are not) capable of.
Figure 1: Our benchmark assesses LLMs’ understanding of symbolic graphics programs in semantic understanding and prediction consistency.
@@
In addition to performing evaluation under the common in-context setting wherein LLMs are used “out-of-the-box” and not finetuned, we also evaluate whether finetuning LLMs on a curated dataset can boost performance. To this end, we propose a new method, Symbolic Instruction Tuning (SIT) to improve the semantic understanding of symbolic programs. The key idea is to collect an instruction-following dataset based on the rendered images. Because the semantic questions of interest are usually easy to answer from the visual input, we take advantage of the rendered images (that correspond to symbolic programs) and query a powerful language-vision model (e.g., GPT-4o) for detailed captioning and question-answering. This leads to a scalable way to collect an instruction-following dataset for symbolic graphics programs. Then, we simply finetune open-source LLMs on this dataset. Our experiments demonstrate that SIT can indeed improve a model’s semantic understanding of symbolic programs. We are excited about future work, motivated by our benchmark, to improve this element of symbolic graphics program understanding. Our contributions are summarized below:
@@
  • We introduce a symbolic MNIST dataset, where the symbolic problem understanding can be extremely challenging so that even the most powerful LLM (i.e., GPT-4o) can only obtain nearly a chance-level performance while the rendered image remains very easy to recognize. This result suggests that how LLMs understand symbolic graphics programs is quite different from human.
  • Finally, we find that SIT can also improve generic instruction following performance.
2 Related Work
Symbolic graphics programs. Generating visual data by procedural modeling with symbolic programs has been essential to computer graphics since its inception, particularly for 2D shapes and 3D geometry. They offer a concise, structured, and interpretable representation, attracting significant attention in the field — refer to [81] for an overview. Common program types include constructive-solid geometry (CSG) [19, 43, 79, 86, 110], computer-aided design (CAD) [28, 48, 49, 85, 104], vector graphics (e.g., SVG) [76, 77], L-systems [31], and customized domains [21, 92, 16, 35, 22, 67]. Among these, SVGs, constructed from primitive shapes like vector paths, curves, or polygons defined by mathematical equations, are widely used in artistic applications such as icon creation and typography. Central to SVGs is the vector path, providing detailed control over graphics and including primitives like rectangles, ellipses, and text. Similarly, procedural 3D geometric modeling, particularly in CAD applications, involves parameterized operations that can be adjusted and re-executed to produce updated geometry. Datasets like the ABC dataset [45] and Fusion 360 Gallery [98] aid CAD research and machine learning, offering hierarchical decomposition, joints, contact surfaces, construction sequences, and shape segmentation based on modeling operations. Our paper focuses on graphics programs of SVG and CAD by introducing a new semantic understanding task that requires a challenging reasoning with the programs.
Relevant benchmarks and datasets. There exist many benchmarks in evaluating different aspects of LLMs: AI safety/ethics [103, 36], out-of-distribution performance [113, 106], API/tool usage [50], code generation [32], etc. Perhaps the most relevant aspect of LLM evaluation to our task is (non-graphics) program understanding abilities, for which a number of papers have explored [93, 42, 32, 63, 55, 62, 87, 33, 9, 13]. As graphics programs can be turned into images via rendering, it is also highly relevant to our tasks to investigate how vision-language models are capable of visual understanding. To evaluate the performance of different models for this related task, numerous types of public multi-modal datasets on generic domains are used, such as human-like image-caption pair datasets (e.g., COCO Caption [10], Nocaps [1], and Flickr30k [72]) and visual question answering datasets (e.g., GQA [38], OK-VQA [68], VQAv2 [30], and Vizwiz [5]). Some specific domain datasets exist as well, such as TextVQA [88] that involves questions on OCR text in images, ScienceQA [83] that involves commonsense questions on scientific topics and Youcook2 [118] that provides temporal information in the form of video clips. These datasets are heterogeneous in both data formats and evaluation metrics and as a result performing a comprehensive evaluation for visual understaning of VLMs is highly tedious. It is also worth pointing out that some more recent datasets like OwlEval [107] and MME [109] are proposed to alleviate this issue with more unified and comprehensive evaluation protocols. Different from all existing benchmarks, SGP-Bench is one of the first benchmarks to evaluate the semantic understanding of symbolic graphics programs.
@@ 3 Semantic Understanding of Symbolic Graphics Programs
We introduce the task of semantic symbolic graphics program understanding. Our goal is to assess to what extent a LLM is able to “understand” a symbolic graphics program, which may begin to belie some latent capability to “visually imagine”. Specifically, we leverage the correspondence between deterministic symbolic graphics programs and rendered images, and then we characterize the understanding of symbolic graphics programs as the semantic understanding of the corresponding rendered image. To do so, we use the performance on question-answering to evaluate the semantic understanding of images. The same set of questions, along with the corresponding symbolic graphics programs, are then used to evaluate the symbolic program understanding of LLMs (the rendered image will not be used here). Figure 2 gives an illustration of symbolic graphics program understanding. The intuition behind this evaluation is that, if an LLM has a good sense of the symbolic graphics and implicit de-rendering, then the LLM should have a rough understanding about its rendered image such that it is able to answer arbitrary semantic questions regarding this rendered image.
@@
Symbolic graphics program understanding can also be viewed as a form of visual question answering in the sense that visual input is represented by a symbolic program representation. Compared to existing vision-language models [54, 53, 119] that encodes images with a text-aligned neural encoder [74], our paper considers the case where the visual input are encoded by a symbolic program that can exactly recover the graphics content. From this perspective, our task aims to study and uncover the potential of using symbolic programs as a representation to perform visual reasoning.
4 SGP-Bench: A Benchmark for Symbolic Graphics Program Understanding
@@ 4.1 Dataset Creation Pipeline
To construct our benchmark, we need questions about a symbolic program based on its rendered image. To build a large benchmark, it is essential to consider how we can effectively scale up the question collection with minimal human efforts. To this end, we use a powerful vision-language model (e.g., GPT-4o) to generate semantic questions based on the rendered images, and then we inspect them manually to make sure that these questions are reasonable and the answer to them is correct. We also run a human study over a randomized set of $500$ of the automatically generated questions along with the corresponding images, and find high agreement (see Appendix B). The overall procedure for our dataset creation is given in Figure 3. In this pipeline, the rendering of symbolic programs and the GPT-4o querying are both scalable and can be done with minimal human involvement. Human annotators then inspect the generated question-answer pairs based on the rendered image, which requires much less efforts than manually writing questions and answers. We emphasize that this program-question-answer triplet creation method is general, as it works for most of the symbolic graphics programs. SVG and 2D CAD programs can directly produce 2D images, so it is straightforward to use this pipeline. For 3D CAD programs, they produce 3D models and we first render them into 2D images with a few fixed camera positions. These rendered images from different views are then combined together to query GPT-4o, and the following procedures are identical to the SVG case.
@@ 4.2 Benchmarking Semantic Understanding
SVG dataset statistics. We collect in total $1,085$ SVG programs covering $19$ categories, and each program has $4$ semantic multiple-choice questions (with $4$ options). We make sure that answers are evenly distributed across $4$ options. Detailed dataset statistics are in Figure 4(a). Our SVG benchmark consists of $5$ different types of questions, including “Semantic”: $1,085$ questions, “Color”: $864$ questions, “Shape”: $1,217$ questions, “Count”: $819$ questions, and “Reasoning”: $355$ questions. “Semantic” tests the global semantic meaning of the object represented by SVG codes, while the other four question types focus on detailed, local understanding of the object. “Color” is color-related questions about specific object parts, which evaluates the localization of the corresponding semantic part. “Count” is about counting the occurrences of certain patterns or semantic parts. “Shape” is about the shape of certain parts of the object, either finding the exact definition of the geometric shape or finding geometric shapes that resemble the object part. Figure 5 gives some SVG examples.
CAD dataset statistics. We collect $2,400$ CAD programs from three different datasets [97, 101, 84]. The CAD dataset consists of $1000$ programs from DeepCAD [101], which forms the 3D subset; $700$ programs from the Fusion360 Reconstruction Dataset [97], which constitutes the 3Drecon. subset; and $700$ programs from SketchGraphs [84], which makes up the 2D subset (as shown in Table 1). Depending on the data, we either render one image or two images from two opposite isometric views (i.e., front and back). Then we feed the images to GPT-4o and generate one semantic multiple-choice question (with 4 options) and its answer. This gives us $2,400$ questions in total. We make sure that ground truth answers are evenly distributed across $4$ options. Detailed dataset statistics are given in Figure 4(b). Some examples from our CAD dataset are provided in Figure 5.
@@
Experimental results and discussion. We find that graphics program “understanding”, as we operationalize it here, is challenging. The average accuracy of all models (proprietary and open-source) is below 65% (ranging from 31% to 63%) on SVG and below 80% (ranging from 29% to 78%) on CAD. Among this, SVG makes it even more difficult for the models to understand as these 2D graphics contain richer semantics. However, the models have substantial performance gains in accordance with the scaling law [111], and this increase, with larger model size, is observed across various open-source LLMs. For example, Llama-3’s score is improved from 0.429 to 0.548 on SVG and from 0.633 to 0.694 when its size increased from 8B to 70B, Qwen-1.5’s from 0.376 to 0.499 on SVG and 0.560 to 0.694 on CAD with the size from 7B to 110B. We also notice consistent improvements given the same model size and model family but from different generations. For example, Qwen-72B from 0.466 to 0.537, Llama-8B from 0.429 to 0.465 and Llama-70B from 0.548 to 0.574. The consistent performance gain on both the SVG and CAD programs indicates that the semantic understanding of symbolic graphics programs, including both 2D semantic understanding and 3D geometry understanding, is a fundamental capability that is aligned with the scaling law of LLMs, even if these models have not been specifically trained for this task.
@@
The semantic understanding of graphics programs can be probed across different aspects, ranging from attribute-level investigations of “color” and “shape” to higher-level discussions of “semantics”, “counting” and “reasoning”. Our benchmark is designed to cover these investigations for LLMs. Most LLMs perform well on color-related questions, followed by shape-related questions, with “count” and “semantic” questions showing progressively lower performance. This consistency is intriguing, as it resembles the coarse-to-fine structure of visual information processing. “Color” is the most visually salient feature, “shape” understanding requires a finer grasp of global and local structures, and “count” and “semantic” questions demand deeper comprehension and knowledge. The difficulty curve is evident, with most open-source models achieving roughly half the accuracy on semantic questions compared to color questions. For instance, the best-performing open-source model, Llama3-405B, achieves 37.6% accuracy on semantics, while reaching 81.6% accuracy on color grounding tasks. While open-source models struggle with “semantic” questions, GPT models perform surprisingly well, with semantics being their second-best category after color grounding.
@@
Avg Semantics Count Color Shape Reason T Avg. SE(2) Avg. T Cons. SE(2) Cons. Avg 3D 3Drecon. 2D
@@
DeepSeek-Coder-V2-16B-Instruct 0.451 0.309 0.379 0.637 0.548 0.268 0.496 0.476 0.902 0.867 0.547 0.611 0.521 0.481
Mistral-Codestral-22b-v0.1 0.491 0.309 0.446 0.698 0.581 0.321 0.503 0.470 0.860 0.816 0.602 0.659 0.577 0.547

### v3

@@ Abstract
Against the backdrop of enthusiasm for large language models (LLMs), there is a growing need to scientifically assess their capabilities and shortcomings. This is nontrivial in part because it is difficult to find tasks which the models have not encountered during training. Utilizing symbolic graphics programs, we propose a domain well-suited to test multiple spatial-semantic reasoning skills of LLMs. Popular in computer graphics, these programs procedurally generate visual data. While LLMs exhibit impressive skills in general program synthesis and analysis, symbolic graphics programs offer a new layer of evaluation: they allow us to test an LLM’s ability to answer semantic questions about the images or 3D geometries without a vision encoder. To semantically understand the symbolic programs, LLMs would need to possess the ability to “imagine” and reason how the corresponding graphics content would look with only the symbolic description of the local curvatures and strokes. We use this task to evaluate LLMs by creating a large benchmark for the semantic visual understanding of symbolic graphics programs, built procedurally with minimal human effort. Particular emphasis is placed on transformations of images that leave the image level semantics invariant while introducing significant changes to the underlying program. We evaluate commercial and open-source LLMs on our benchmark to assess their ability to reason about visual output of programs, finding that LLMs considered stronger at reasoning generally perform better. Lastly, we introduce a novel method to improve this ability – Symbolic Instruction Tuning (SIT), in which the LLM is finetuned with pre-collected instruction data on symbolic graphics programs. Interestingly, we find that SIT not only improves LLM’s understanding on symbolic programs, but it also improves general reasoning ability on various other benchmarks.
[…]
@@ 1 Introduction
What are large language models (LLMs) capable of? Recent studies [5, 58] have shown that LLMs are able to generate generic computer programs, indicating a degree of pragmatic understanding of the symbolic structure of programs. Motivated by this progress, we focus on another important family of computer programs, called symbolic graphics programs, where a graphics content (e.g., image, 3D asset) can be generated by running a program. We are interested in the following question: Can large language models “understand” symbolic graphics programs?
Before trying to answer this question, we start by defining what we consider “understanding” of symbolic graphics programs, in the context of this work. Because a (deterministic) graphics program can be uniquely rendered to an image (the graphics programs we consider here), we characterize LLMs’ understanding of the graphics program as the semantic understanding of the corresponding rendered image. More specifically, we approximate such a semantic visual understanding by the ability to correctly answer semantic questions only based on the raw program input. These semantic questions are generated based on the rendered image, such that they are easy to answer given the image and yet challenging given only the program as text prompt. Guided by this insight, we propose a generic pipeline for creating benchmarks that can evaluate this particular ability of LLMs to understand symbolic graphics programs, while requiring minimal human effort. While we of course recognize that there are other elements of visual reasoning that characterize understanding in humans, and that we ought to evaluate in machine intelligence, we believe that our benchmark provides insight into one element of “understanding” of symbolic graphics programs that helps assay what current LLMs are (and are not) capable of.
Figure 1: Our benchmark assesses LLMs’ understanding of symbolic graphics programs in semantic understanding and prediction consistency. Note that the LLM can only see symbolic graphics programs and the corresponding questions. The rendered images are not input to the LLM.
@@
In addition to performing evaluation under the common in-context setting wherein LLMs are used “out-of-the-box” and not finetuned, we also evaluate whether finetuning LLMs on a curated dataset can boost performance. To this end, we propose Symbolic Instruction Tuning (SIT). The key idea is to collect an instruction dataset based on the rendered images. Because the semantic questions of interest are usually easy to answer from the visual input, we take advantage of the rendered images (that correspond to symbolic programs) and query a powerful language-vision model (e.g., GPT-4o) for detailed captioning. This leads to a scalable way to collect an instruction dataset for symbolic graphics programs. Then, we simply finetune open-source LLMs on this dataset. Our experiments demonstrate that SIT can improve a model’s semantic understanding of symbolic programs, and more importantly, its general reasoning ability. Our contributions are summarized below:
@@
  • We introduce a symbolic MNIST dataset, where the symbolic program understanding can be challenging so that even a powerful LLM (i.e., GPT-4o) can only obtain nearly a chance-level performance while the rendered image remains very easy to recognize.
  • Finally, we find that SIT can also improve general reasoning performance.
@@ 2 Semantic Understanding of Symbolic Graphics Programs
We introduce the task of semantic symbolic graphics program understanding. Our goal is to assess to what extent a LLM is able to “understand” a symbolic graphics program, which may begin to belie some latent capability to “visually imagine”. Specifically, we leverage the correspondence between deterministic symbolic graphics programs and rendered images, and then we characterize the understanding of symbolic graphics programs as the semantic understanding of the corresponding rendered image. To do so, we use the performance on question-answering to evaluate the semantic understanding of images. The same set of questions, along with the corresponding symbolic graphics programs, are then used to evaluate the symbolic program understanding of LLMs (the rendered image will not be used here). Figure 2 gives an illustration of symbolic graphics program understanding. The intuition behind this evaluation is that, if an LLM has a good sense of the symbolic graphics and implicit de-rendering, then the LLM should have a rough understanding about its rendered image such that it is able to answer arbitrary semantic questions about the rendered image.
Figure 3: A qualitative example of CAD reasoning.
@@
Symbolic graphics program understanding can be viewed as a form of visual question answering in the sense that visual input is represented by a symbolic program representation. Compared to current vision-language models [56, 55, 121] that encodes images with a text-aligned encoder [76], we consider the case where the visual input are encoded by a symbolic program that can exactly recover the graphics content. From this perspective, our task aims to uncover the potential of using symbolic programs as a representation to perform visual reasoning.
3 Why is Understanding Symbolic
Graphics Programs Interesting?
To showcase why semantic understanding of symbolic graphics problems require multiple sophisticated reasoning abilities, we provide a few qualitative examples of LLM’s output in Figure 3 and Figure 4. Figure 3 shows a qualitative example of how OpenAI-o1 reasons over the CAD program. The reasoning process is highly nontrivial, as it requires multiple reasoning abilities, such as numeric perception, spatial reasoning, geometric understanding, long-range planning and common sense. This is also part of the reason that SGP-Bench can well evaluate the general reasoning ability of LLMs. In Figure 4, we query different LLMs (from weak to strong: Llama-3.1-8B, Llama-3.1-70B, OpenAI-o1) by asking which digit this given SVG program represents. We can observe that all LLMs start to reason from the low-level curves and gradually build up its understanding from local components to semantic elements. Specifically, we can observe that LLMs understand the symbolic graphics program through line-by-line reasoning and then combine the detailed reasoning into an overall semantic understanding. This process is particularly intriguing because it shows that LLMs understand symbolic programs through actual reasoning rather than memorization. More interestingly, the more powerful LLMs, e.g., OpenAI-o1, show better general understanding of the symbolic graphics program and fine-grained grounding of the program, which is consistent with the results on SGP-Bench.
Figure 4: Qualitative examples of how LLMs reason over the symbolic program and obtain their answers.
Figure 5: OpenAI-o1 still suffers from the spurious correlation from the Ebbinghaus illusion while reasoning over images (a). In contrast, OpenAI-o1 works perfectly fine while reasoning over symbolic graphics programs directly (b) or indirectly (c).
To highlight the importance of symbolic graphics programs, we give another example in Figure 5, where a powerful LLM like OpenAI-o1 can suffer from the visual spurious correlation, and in contrast, this spurious correlation can be avoided when using the symbolic graphics program as the visual representation. Specifically, we construct a visual example that resembles the Ebbinghaus illusion, but the conclusion differs from the classic Ebbinghaus illusion (i.e., two orange circles are of the same size, although they look different) as we intentionally make one of the orange circle obviously larger than the other (shown in Figure 5). Then we feed this curated example to OpenAI-o1 and ask which orange circle is larger. We compare three cases: (a) image input; (b) symbolic program input; and (c) image input with indirect symbolic program. As long as we first ask the LLM to translate the image into symbolic program, the LLM will no longer suffer from the spurious correlation. The same phenomenon also appears for the Müller-Lyer illusion using Claude-3.5-Sonnet (Figure 6). These examples validate the advantages of reasoning the visual world through symbolic graphics programs.
Figure 6: Claude-3.5-Sonnet also suffers from the spurious correlation from the Müller-Lyer illusion.
4 A Benchmark for Symbolic Graphics Program Understanding
@@ 4.1 Dataset Creation Pipeline
To construct our benchmark, we need questions about a symbolic program, based on its rendered image. To build a large benchmark, it is essential to consider how we can scale up the question collection effectively, with minimal human effort. To this end, we use a powerful vision-language model (e.g., GPT-4o) to generate semantic questions based on the rendered images, and then we inspect them manually to make sure that these questions are reasonable and the answer to them is correct. We also run a human study over a randomized set of $500$ of the automatically generated questions along with the corresponding images, and find high agreement (see Appendix B). The overall procedure for our dataset creation is given in Figure 7. In this pipeline, the rendering of symbolic programs and the GPT-4o querying are both scalable and can be done with minimal human involvement. Human annotators then inspect the generated question-answer pairs based on the rendered image, which requires much less efforts than manually writing questions and answers. We emphasize that this program-question-answer triplet creation method is general, as it works for most of the symbolic graphics programs. SVG and 2D CAD programs can directly produce 2D images, so it is straightforward to use this pipeline. For 3D CAD programs, they produce 3D models and we first render them into 2D images with a few fixed camera positions. These rendered images are used to query GPT-4o, and the following procedures are identical to the SVG case.
@@ 4.2 Benchmarking Semantic Understanding
SVG dataset statistics. We collect in total $1,085$ SVG programs covering $19$ categories, and each program has $4$ semantic multiple-choice questions (with $4$ options), resulting in a total of $4,340$ questions. We make sure that answers are evenly distributed across $4$ options. Detailed dataset statistics are in Figure 8(a). Our SVG benchmark consists of $5$ different types of questions, including “Semantic”: $1,085$ questions, “Color”: $864$ questions, “Shape”: $1,217$ questions, “Count”: $819$ questions, and “Reasoning”: $355$ questions. “Semantic” tests the global semantic meaning of the object represented by SVG codes, while the other four question types focus on detailed, local understanding of the object. “Color” is color-related questions about specific object parts, which evaluates the localization of the corresponding semantic part. “Count” is about counting the occurrences of certain patterns or semantic parts. “Shape” is about the shape of certain parts of the object, either finding the exact definition of the geometric shape or finding geometric shapes that resemble the object part. Figure 9 gives some SVG examples.
CAD dataset statistics. We collect $2,400$ CAD programs from three different datasets [101, 105, 86]. The CAD dataset consists of $1000$ programs from DeepCAD [105], which forms the 3D subset; $700$ programs from the Fusion360 Reconstruction Dataset [101], which constitutes the 3Dcomplex subset; and $700$ programs from SketchGraphs [86], which makes up the 2D subset (as shown in Table 1). Different from SVG, there is no generally established syntax for building CAD models from graphics codes; each of our 3 CAD subsets follows a different language syntax, with varying levels of complexity. When benchmarking LLMs for CAD tasks, we include domain-specific language syntax rules as part of the input prompt. The LLM is required to apply in-context learning of this syntax and answer the test questions. Then we feed the renderings to GPT-4o and generate one semantic multiple-choice question (with 4 options) and its answer. This gives us $2,400$ questions in total. We make sure that ground truth answers are evenly distributed across $4$ options. Detailed dataset statistics are given in Figure 8(b). Some examples from our CAD dataset are provided in Figure 9.
@@
Experimental results and discussion. We find that graphics program understanding, as we operationalize it here, is challenging. The average accuracy of all models (proprietary and open-source) is below 70% (ranging from 30% to 67%) on SVG and below 75% (ranging from 28% to 74%) on CAD. Among these, SVG makes it more difficult for the models to understand as these 2D graphics contain richer semantics. Significant performance improvements are observed in line with scaling laws [113], as larger model sizes consistently lead to gains across various open-source LLMs. For example, Llama-3’s score is improved from 0.429 to 0.548 on SVG and from 0.633 to 0.694 when its size increased from 8B to 70B, Qwen-1.5’s from 0.376 to 0.499 on SVG and 0.486 to 0.632 on CAD with the size from 7B to 110B. We also notice consistent improvements given the same model size and model family but from different generations. For example, Qwen-72B from 0.466 to 0.537, Llama-8B from 0.429 to 0.465 and Llama-70B from 0.548 to 0.574. The consistent performance gain on both the SVG and CAD indicates that the semantic understanding of symbolic graphics programs, including both 2D and 3D understanding, is a fundamental capability that is aligned with the scaling law of LLMs, even if these models have not been specifically trained for this task.
@@
The semantic understanding of graphics programs can be probed across different aspects, ranging from attribute-level investigations of “color” and “shape” to higher-level discussions of “semantics”, “counting” and “reasoning”. Our benchmark is designed to cover these investigations for LLMs. Most LLMs perform well on color-related questions, followed by shape-related questions, with “count” and “semantic” questions showing progressively lower performance. This consistency is intriguing, as it resembles the coarse-to-fine structure of visual information processing. “Color” is the most visually salient feature, “shape” understanding requires a finer grasp of global and local structures, and “count” and “semantic” questions demand deeper comprehension and knowledge. The difficulty curve is evident, with most open-source models achieving roughly half the accuracy on semantic questions compared to color questions. For instance, the best-performing open-source model, Llama3.1-405B, achieves 37.6% accuracy on semantics and 81.6% accuracy on color grounding tasks. While open-source models struggle with “semantic” questions, ChatGPT performs quite well, with semantics being their second-best category after color grounding.
@@
Avg Semantics Count Color Shape Reason T Avg. SE(2) Avg. T Cons. SE(2) Cons. Avg 3D 3Dcomplex 2D
@@
DeepSeek-Coder-V2-16B 0.451 0.309 0.379 0.637 0.548 0.268 0.496 0.476 0.902 0.867 0.547 0.611 0.521 0.481
Codestral-22B-v0.1 0.491 0.309 0.446 0.698 0.581 0.321 0.503 0.470 0.860 0.816 0.602 0.659 0.577 0.547